OpenStudy (anonymous):

Find the slope of the line passing through the points (3, 4) and (2, 1). (Points : 1) Option A: 1 Option B: -1 Option C: 3/5 Option D: -5/3

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

I don't quite understand......

hartnn (hartnn):

there is formula above... y1-y2 = 4-1 = ... ?

OpenStudy (anonymous):

I think I get it now. Thanx ^_^

hartnn (hartnn):

what u got your final slope as ? welcome ^_^